OpenStudy (anonymous):

Find the distance CB for the points C(8, –6) and B(6, 7). ( answer more appreciated, im on e2020)

OpenStudy (johnweldon1993):

Distance between 2 points \[\large \sqrt{(x_2 - x_1)^2 + (y_2 - y_1)^2}\] You have your 2 points...so C(8, –6) (x1 , y1) and B(6, 7) (x2 , y2) Just plug in those points and solve for distance

OpenStudy (johnweldon1993):

So in your case \[\large \sqrt{(6 - 8)^2 + (7 + 6)^2}\] When simplified down that is \[\large \sqrt{(-2)^2 + (13)^2}\] \[\large \sqrt{(4 + 169)}\] \[\large \sqrt{(173)}\] would be the distance...that is exact form...if you want the decimal form...just do the square root of 173

OpenStudy (anonymous):

No that's what I needed thank you

OpenStudy (johnweldon1993):

Anytime :)